40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jan 14 01:55:26 2011, kiwamu, Update, LSC, X arm locked with C1LSC digital control 
    Reply  Fri Jan 14 11:29:00 2011, kiwamu, Update, LSC, expected open loop TF of X arm locking xarm_oltf.png
    Reply  Fri Jan 14 12:34:08 2011, Koji, Update, LSC, X arm locked with C1LSC digital control 
Message ID: 4153     Entry time: Fri Jan 14 01:55:26 2011     Reply to this: 4154   4156
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: X arm locked with C1LSC digital control 

 [Koji, Kiwamu]

 We succeeded in locking the X arm with the C1LSC digital control.

As we did on the day before yesterday, the feedback signal goes to MCL (#4141), but this time the signal is transfered from C1LSC through the RFM.

 


 (key points) 

- checking the state of the analog whitening filters at C1LSC rack.

   We took the transfer function of them and found that they were always on regardless of the clicking any buttons on medm.

To cancel the filter shape of the whitening, we put an unWhitening filter so that these transfer functions becomes flat in total.

The whitening filter approximately has : pole:150Hz, pole:150Hz, zero:15Hz, zero:15Hz (although these numbers came from by our eye ball fitting)   

 

 - demodulation phase adjustment

   We performed the same measurement as that of Suresh and Koji did yesterday (#4143) to adjust the phase of the PDH demodulation.

By changing the cable length we roughly adjusted the I-phase to eventually ~10 deg, which is close enough to 0 deg.

(probably some more efforts should be made as a part of daytime tasks)

Note that we are currently using the REFL33 demodulation board for this purpose (#4144). The LO power we put is about 16dBm.

The angle between I and Q at 11MHz is actually almost 90 deg.

This fact has been confirmed by putting a sinusoidal signal with a slightly different frequency (~100Hz) from that of the LO onto the RF input.

 

 - attenuation of RF signal

  Since the PDH signal taken by C1LSC's ADC had been saturated somewhat, we introduced a ND filter of 10 on the photo diode to attenuate the RF signal.

As a result the amplitude of the PDH signal on dataviewer became more reasonable. No more saturations.

 

(some notes)

 unWhitening filter           pole:15Hz. pole:15Hz, zero:150Hz, zero:150Hz

 C1LSC_MC_FM1   pole:1kHz, zero:10Hz

 Gain in digital control       G ~ -1

measured UGF  ~  200-300 Hz

 measured RFM delay ~ 125 usec 

ELOG V3.1.3-